# Decoding News Narratives: A Critical Analysis of Large Language Models in Framing Bias Detection

### v2

@@ 3.2.2 Few-Shot Prompting
To understand how the number of examples affects the model’s accuracy, we evaluate few-shot models in two different configurations. The first configuration includes a minimal set of two examples, one framed and one not framed. This setup helps us measure the baseline impact of including examples on model performance. The second configuration includes eight examples, from which four examples are framed headlines.
@@
We assess four scenarios to determine how the relevance of examples to the headline topics affects the model’s ability to detect framing:
@@
Varied in-Domain: In this setting framed examples cover four diverse aspects of gun violence777I.e., politics, public/school safety, race/ethnicity, and social/culture. to test the model’s adaptability to a range of in-domain cues.
Cross-Domain: To evaluate the model’s generalisation skills across topics, we use examples from completely different domains, such as immigration, in the few-shot prompts.
@@ 3.2.3 Explainable Prompting
| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | ||
| Zero-Shot | No Definition | 0.46 | 49.06 | 17.96 | 49.21 | 0 | 0 | 25.86 | 52.74 | 34.33 | 46.19 | 51.01 | 49.06 |
| +Definition | 0.15 | 50.13 | 1.21 | 50.10 | 0 | 0 | 0.61 | 50.06 | 42.27 | 48.06 | 37.63 | 43.96 | |
| Few-Shot | 2 examples | 0.31 | 50.06 | 0.31 | 50.06 | 0 | 0 | 3.12 | 49.98 | 42.27 | 48.06 | 37.63 | 43.96 |
| 8 (focused in-domain) | 0.15 | 50.17 | 3.84 | 50.10 | 0 | 0 | 4.83 | 50.17 | 16.78 | 47.53 | 36.85 | 44.69 | |
| 8 (varied in-domain) | 0 | 0 | 0.15 | 49.94 | 0 | 0 | 1.66 | 49.94 | 22.78 | 45.96 | 35.03 | 45.27 | |
| 8 (cross domain) | 5.30 | 49.37 | 2.95 | 49.60 | 0 | 0 | 4.99 | 50.33 | 40.23 | 45.11 | 51.64 | 46.88 | |
| 8 (mixed domain) | 0.31 | 50.21 | 1.51 | 50.02 | 0 | 0 | 3.84 | 50.13 | 34.39 | 44.58 | 44.05 | 43.43 | |
4 LLMs’ Reliability in Detecting Framing Bias
The evaluation results for GPT and FLAN-T5 models across different settings are presented in Table 1 and Table 2, respectively. We can summarise the notable findings of the results of as follows:
Explainable Prompting Enhances Reliability:
Explainable prompting consistently yielded more reliable outcomes in both zero- and few-shot variations, as shown by the reduced variance in accuracy and F scores. This underscores the importance of explainable settings for social science research on framing bias.
Optimal Performance in Few-Shot with Diverse Examples:
The GPT-4 model achieved the highest accuracy and F scores in few-shot scenarios with a wide range of in-domain examples (8 varied in-domain). This finding suggests that incorporating diverse, relevant in-domain examples significantly improves model performance when the news headline domain is known.
Tendency Towards Framing Bias Classification:
In explainable settings, both GPT-3.5 and GPT-4 models exhibit a higher tendency to classify headlines as framed. This is indicated by higher F scores for framed headline detection alongside lower overall accuracy, suggesting a bias towards classifying headlines as framed when prompted to identify framing.
Challenges with Cross-Domain Examples:
F scores dropped in cross-domain settings without explainable prompts, falling below those in zero-shot settings with definitions. This highlights the challenges of applying advanced models to new or mixed domains, where headline topics vary widely. Therefore, for new domain applications, we suggest to either employ a zero-shot setting with task definition, or opting for explainable prompts when examples are used in the input prompts. However, for large-scale data analysis, opting for a zero-shot approach without explanations could be more efficient, albeit at the potential cost of reduced reliability compared to the explainable setting.
Inherent Challenge in Framing Bias Detection:
Across both models, the highest F scores observed were in the range of 60-70 points, highlighting the inherent difficulty of detecting framing bias. Framing detection requires understanding subtle language nuances, contextual cues, and the intended message framing, which are complex and context-dependent tasks. Furthermore, the performance between GPT-3.5 and GPT-4 did not significantly differ, particularly in more realistic use cases like zero-shot and cross-domain settings. Considering the higher cost of using GPT-4’s API, GPT-3.5 might be a more cost-effective choice for framing bias detection in large scale analysis.
FLAN-T5 Performance Lag:
The results from Table 2 show that FLAN-T5 models significantly underperform compared to GPT models, with the best F score reaching only 51 points. This indicates that smaller models like FLAN-T5 may not yet be viable for complex tasks such as framing bias detection without additional task-specific fine-tuning or more extensive training data.
5 Analysis
@@
| Headline: Pregnant woman who killed intruder in justified shooting now faces felony gun charge due to previous marijuana conviction. |
@@
| GPT-4 Explanation: The claim is framed. It selectively emphasizes on the individual’s status as a pregnant woman and the nature of the shooting |
| as justified, while sidelining the legal issue of her previous marijuana conviction which is directly linked to the current felony charge. |
| The framing encourages empathy for the person involved by prominently highlighting her pregnancy and the self-defense aspect, provoking an emotional response. |
| GPT-3.5 Explanation: This claim appears to accurately represent the situation. It states that a pregnant woman killed an intruder in a justified |
| shooting and is now facing a felony gun charge due to a previous marijuana conviction. The claim does not seem to contain any misleading or false information. |
@@
The Impact of Emotional Language on Framing Bias Detection
Our analysis of GPT-4’s errors reveals a consistent pattern: the model frequently interprets emotional language as an indicator of framing bias, irrespective of the headline’s actual framing. This tendency likely stems from the observation that emotionally charged language often accompanies framed headlines, evoking specific frames and eliciting emotional responses from readers, thereby influencing their perception of the news narrative Valkenburg et al. (1999). However, this leads to a biased error in the model, mistakenly identifying expressions of emotion as indicative of framing bias. Such a pattern highlights the challenge of distinguishing between genuine emotional expression and intentional framing bias. Making this distinction is essential for enhancing the model’s precision in analysing news narratives. Table 3 provides examples of such misclassifications by GPT-4 in the explainable zero-shot setting. On the other hand, GPT-3.5 seems to be less affected by the emotional language.
| GPT-3.5 | GPT-4 | Flan-T5 | ||||||||
| Small | Base | Large | ||||||||
| F | Acc. | F | Acc. | F | Acc. | F | Acc. | F | Acc. | |
| Clear | 71.43 | 70.15 | 75.81 | 77.61 | 3.23 | 42.86 | 9.23 | 43.81 | 49.6 | 40.00 |
| Contested | 41.43 | 38.81 | 3.60 | 20.15 | 3.03 | 39.62 | 0 | 0 | 65.60 | 59.43 |
| Clear | Contested | |
| Agreement | 69.40 | 58.95 |
| Agreement GL | 58.20 | 16.45 |
| GPT-3.5 | GPT-4 | |||||||
| Explainable | Explainable | |||||||
| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
| Framed | Not-Framed | |
| GPT-3.5 Turbo | 84 | 38 |
| GPT-4 | 108 | 22 |
| Flan-T5 Small | 1 | 129 |
| Flan-T5 Base | 0 | 130 |
| Flan-T5 Large | 32 | 98 |
Clear vs. Contested Cases of Framing Annotations
During our initial analysis, we observed discrepancies in some data annotations, with some annotations being potentially incorrect. Additionally, there were ambiguous cases where multiple labels could be justified based on different interpretations. In this regard, we manually examined half of the headlines in GVFC, i.e., 1300 headlines. Within this subset, we identified 134 contested annotations. Examples include “Live: Trump visits Pittsburgh after synagogue shooting” and “Shopify bans sale of certain firearms, accessories”, both annotated as framed. However, not all 134 cases are necessarily wrong annotations. Some of them, such as “Thousands gather to honor victims of the mass shooting with tears, candlelight, and song”, present more nuanced challenges in framing detection.88863.4% of headlines in this subset are annotated as framed in the dataset.
Additionally, we selected another set of 134 headlines where the presence or absence of framing bias was more apparent. For instance, headlines such as “Two dead including shooter at Florida yoga studio” and “Parkland school shooter blames massacre on a ’demon’ voice” serve as clear examples of non-framed and framed headlines that are also annotated correctly, respectively. We call this subset the “Clear” subset.
Table 4 reports the results of evaluating our models on these two distinct subsets. As expected, both GPT models exhibited strong performance on the Clear subset. However, they have a considerably low results on the contested subset, with GPT-4’s F score dropping to 3.6 points.
We have also calculated how often GPT-4 and GPT-3.5’s predictions aligned within these subsets. These agreement ratios are shown in Table 5. We observe that while the models reached an agreement of about 59% on the contested subset, only 16% of these agreed-upon predictions matched the gold labels. This suggests that these models’ can be useful in flagging potential annotation inaccuracies within existing or new datasets. As an example, both models identify the “ Muslim Americans raise more than $200,000 for those affected by Pittsburgh synagogue shooting” as not framed while it is annotated as framed in the data.
Impact of Prompt Variability
One of the challenges with prompt-based techniques is their sensitivity to the wording of input prompts; even minor variations can lead to drastically different performances. We examined this variability by using various input examples in focused in-domain, varied in-domain, cross-domain, and mixed-domain settings as shown in Table 1. To further examine the impact of prompt wording on framing bias detection results, we altered the prompt from “Decide whether this claim is framed or not” to “Decide whether this claim contains framing bias” in our zero-shot experiments. Table 6 shows these findings. Our result indicates that the explainable setting yields the most consistent results, where different formulations of the question minimally affects outcome variability. On the other hand, we observe the most significant performance change in the GPT-4 model without a defined task or explanation, where the F score drops from 64 to 47 points. Given that the explainable setting requires more computational resources, including a task definition is a more efficient alternative for decreasing variability in results, and therefore enhancing their reliability.
Evaluation in the Wild
Until now, our results and analysis were centered on the GVFC dataset, which is limited to the narrow subject of gun violence within the U.S. To assess the models in a scenario closer to real-world conditions, where the topics of headlines are varied and unknown, we collected a new dataset from a website known for its emphasis on news framing.999https://newsframes.wordpress.com/category/headlines/ This collection consists of 130 headlines spanning diverse subjects such as British weather, health, and European matters.
We assessed the models on this novel dataset employing the explainable few-shot setting with 8 cross-domain examples, which resulted in reliable results in Section 4. Table 7 shows the counts of headlines predicted as framed or not framed by each model. Given the dataset’s origin from a website known for featuring framed headlines, we anticipated the majority of headlines to be framed. We observe that the GPT-4 model predominantly identified headlines as framed. However, considering GPT-4’s tendency towards classifying inputs as framed, a promising avenue for future research involves developing a dataset specifically aimed at challenging the models with not-framed cases. This could include not-framed headlines that still incorporate emotional language, as discussed in the beginning of this section.
6 Conclusions
This work advances our understanding of the performance of LLMs such as GPT-3.5 Turbo, GPT-4, and Flan-T5 (small, base, and large) in detecting framing bias within news headlines. Through a comprehensive experimental approach that included zero-shot, few-shot, and explainable prompting settings, we found that GPT-4 excels in few-shot scenarios with a variety of in-domain examples, highlighting its superior capability in recognising framing bias. However, GPT-4 also displayed a tendency to classify headlines as framed, suggesting a valuable line of research could involve testing these models on datasets containing non-framed headlines that potentially also incorporate emotional language, identified as another potential weakness. In fact, our research explores the impact of emotional language on framing bias detection, with GPT-3.5 showing resilience against emotional language.
The study also identifies a performance shortfall in Flan-T5, underscoring the challenges smaller models face in complex detection tasks without specific fine-tuning. Our findings indicate that F1 scores drop in cross-domain scenarios with non-explainable prompts, suggesting the effectiveness of zero-shot approaches with clear definitions or explainable prompts for new domain applications. The study further demonstrates the utility of these models in identifying potential annotation inaccuracies in new or already existing datasets.
7 Limitations
The findings of this study have to be seen in light of some limitations. For example, our evaluation focuses solely on English-language content, leaving space for further investigation on other languages to explore our findings’ applicability to non-English contexts. This limitation suggests a need for further investigation into the performance of LLMs across different languages and cultural contexts to fully assess the potential use of these models in social science research for detecting framing bias and analysing media narratives.
8 Acknowledgements
This work was supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation [grant number EP/S023062/1].

### v3

@@ 3.4 Few-Shot Prompting
To understand how the number of examples affects the model’s accuracy, we evaluate few-shot models in two different configurations. The first configuration includes a minimal set of two GVFC examples, one framed and one not framed. This setup helps us measure the baseline impact of including examples on model performance. The second configuration includes eight examples, from which four examples are framed headlines.
@@
We assess four scenarios to determine how the relevance of the examples to the headline topics in the test set, if known a priori, affects the model’s ability to detect framing:
@@
Varied in-Domain: In this setting framed examples cover four diverse aspects of gun violence666I.e., politics, public/school safety, race/ethnicity, and social/culture. to test the model’s adaptability to a range of in-domain cues.
Cross-Domain: To evaluate the model’s performance on topics not known in advance, we use examples from completely different domains, such as immigration.
@@ 3.5 Explainable Prompting
4 Evaluating Model Reliability in Detecting Framing
The evaluation results for the GPT models across different settings are presented in Table 1, while the results for the FLAN-T5 models can be found in Appendix A. We can summarise the notable findings of the results as follows:
Explainable Prompting Enhances Reliability:
Explainable prompting consistently yielded more reliable outcomes in both zero- and few-shot variations, as shown by the reduced variance in accuracy and F1 scores across different prompting settings.
Optimal Performance in Few-Shot with Diverse Examples and Explanation:
The GPT-4 model achieved the highest accuracy and F1 scores in few-shot scenarios with a wide range of in-domain examples (8 varied in-domain). This result suggests that when the topic of news headlines is known in advance, incorporating this information into few-shot examples can considerably improve performance.
Challenges with Cross-Domain Examples:
F1 scores dropped in cross-domain settings without explainable prompts, falling below those in zero-shot settings with definitions. This highlights the challenges of applying advanced models to new or mixed domains, where headline topics vary widely. Therefore, for new domain applications, we suggest to either employ a zero-shot setting with task definition, or opting for explainable prompts when examples are used in the input prompts. We test our suggested settings on real world conditions in our analysis in section 5.
Inherent Challenge in Framing Detection:
Across both models, the highest F1 scores observed were in the range of 60-70%, highlighting the inherent difficulty of detecting framing. Framing detection requires understanding subtle language nuances, contextual cues, and the intended message framing, which are complex and context-dependent tasks. Furthermore, the performance between GPT-3.5 and GPT-4 did not significantly differ, particularly in more realistic use cases like zero-shot and cross-domain settings. Considering the higher cost of using GPT-4’s API, GPT-3.5 might be a more cost-effective choice for framing detection in large scale analysis.
FLAN-T5 Performance Lag:
Out-of-the-box FLAN-T5 models significantly underperform compared to GPT models, with the best F1 score reaching only 51%, which is negligibly above the majority baseline. The detailed results can be found in Appendix A. This performance lag indicates that smaller models like FLAN-T5 may not yet be viable for complex tasks such as framing detection without additional task-specific fine-tuning or more extensive training data. To assess this, we have fine-tuned the best performing model (FLAN-T5 Large) on the GVFC and tested it. This analysis is discussed in section 5.
5 Analysis
The Impact of Emotional Language on Framing Detection
Our analysis of GPT-4’s errors reveals a consistent pattern: the model frequently interprets emotional language as an indicator of framing, irrespective of whether the headline is actually framed or not. This tendency likely stems from the observation that emotionally charged language often accompanies framed headlines, evoking specific frames and eliciting emotional responses from readers, thereby influencing their perception of the news narrative Valkenburg et al. (1999).
However, this leads to a biased error in the model, mistakenly identifying expressions of emotion as indicative of framing. Our findings are consistent with recent research on the weight of emotional charge in sentiment classification tasks with LLMs Burnham (2024). Such a pattern highlights the challenge of distinguishing between genuine emotional expression and intentional framing. Making this distinction is essential for enhancing the model’s precision in analysing news narratives. Appendix C provides examples of such misclassifications by GPT-4 in the explainable zero-shot setting. On the other hand, GPT-3.5 seems to be less affected by emotional language.
| Flan-T5 | GPT-3.5 | GPT-4 | ||||||||
| Small | Base | Large | ||||||||
| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
| Clear | 2.70 | 44.62 | 10.39 | 46.92 | 51.61 | 42.31 | 71.43 | 70.15 | 75.81 | 77.61 |
| Contested | 2.33 | 36.36 | 0 | 34.85 | 67.47 | 59.09 | 41.43 | 38.81 | 3.60 | 20.15 |
Clear vs. Contested Cases of Framing Annotations
During our initial analysis, we observed discrepancies in the GVFC data annotations, with some annotations being potentially incorrect. Additionally, there were ambiguous cases where multiple labels could be justified based on different interpretations. In this regard, we manually examined half of the headlines in GVFC, i.e., 1300 headlines. Within this subset, we identified 134 contested annotations. Examples include “Live: Trump visits Pittsburgh after synagogue shooting” and “Shopify bans sale of certain firearms, accessories”, both annotated as framed. However, not all 134 cases are necessarily wrong annotations. Some of them, such as “Thousands gather to honor victims of the mass shooting with tears, candlelight, and song”, present more nuanced challenges in framing detection. In this dataset, 63.4% of headlines in this subset are annotated as framed.
Additionally, we selected another set of 134 headlines where the presence or absence of framing was more apparent. For instance, headlines such as “Two dead including shooter at Florida yoga studio” and “Parkland school shooter blames massacre on a ’demon’ voice” serve as clear examples of non-framed and framed headlines that are also annotated correctly, respectively. We call this subset the “Clear” subset.
Table 2 reports the results of evaluating our models on these two distinct subsets. As expected, both GPT models exhibited strong performance on the Clear subset. However, they have a considerably low results on the contested subset, with GPT-4’s F1 score dropping to 3.60%.
We have also calculated how often GPT-4 and GPT-3.5’s predictions aligned within these subsets. These agreement ratios are shown in Table 3. We observe that while the models reached an agreement of about 59% on the contested subset, only 16% of these agreed-upon predictions matched the gold labels. Having manually checked the annotations for the contested subset, our finding suggests that these models’ could be useful in flagging potential annotation inaccuracies within existing or new datasets, primarily when there is agreement between them. As an example, both models identify the “ Muslim Americans raise more than $200,000 for those affected by Pittsburgh synagogue shooting” as not framed while it is annotated as framed in the data.
| Clear | Contested | |
| Agreement | 69.40 | 58.95 |
| Agreement GL | 58.20 | 16.45 |
| GPT-3.5 Turbo | GPT-4 | |||||||
| Explainable | Explainable | |||||||
| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | |
Impact of Prompt Variability
One of the challenges with prompt-based techniques is their sensitivity to the wording of input prompts; even minor variations can lead to drastically different performances. We examined this variability by using various input examples in focused in-domain, varied in-domain, cross-domain, and mixed-domain settings as shown in Table 1. To further examine the impact of prompt wording on framing detection results, we altered the prompt from “Decide whether this claim is framed” to “Decide whether this claim contains framing bias” in our zero-shot experiments.
Table 4 shows these findings. The results show that even a simple change in the input prompt can result in a notable performance difference. For instance, there was a 15-point drop in F1 score for GPT-4 when the prompt did not include a definition of framing. However, the performance variance is considerably lower when the prompts include a definition of framing.
Evaluation In The Wild (ITW) Dataset
Until now, our results and analysis were centered on the GVFC dataset, which is limited to the narrow subject of gun violence within the U.S. To assess the models in a scenario closer to real-world conditions, where the topics of headlines are varied and unknown, we collected a set of framed headlines from a website known for its emphasis on news framing777https://newsframes.wordpress.com/category/headlines/. To make the dataset more balanced, we have also collected instances of not framed news headlines from newspapers such as BBC, The Guardian and Daily Mail. In this dataset, an example of framed headline is “EU plot to scrap Britain”, while an example of not framed headline is “Around 80 migrants including children rescued from Channel”.
The entire dataset consists of 157 headlines spanning diverse subjects such as British weather, health, migration and European matters, and it contains 83 framed headlines and 74 not framed headlines, with a Majority Baseline Accuracy of 52.86%. The annotations were confirmed by one of the authors, who is an expert in framing.
We assessed the models on this novel dataset employing both zero and few-shot settings. For the zero-shot, motivated by our results in the different wording evaluation (Table 4), we used both the standard and explainable settings of the Zero-Shot + Definition prompt. For the few-shot settings with 8 different examples, we used the examples employed in the varied in-domain and cross-domain settings of the Experiments in Section 4. Table 5 shows the results of our In the Wild evaluation, where GPT-4 reached 80.25% accuracy in one of the Few-Shot settings. However, GPT-4 remains sensitive to the examples provided in the prompt, as shown by the 6.6 points difference between the two Few-Shot settings.
| GPT-3.5 Turbo | GPT-4 | |||
| F1 | Acc. | F1 | Acc. | |
| ZS + definition (S) | 74.16 | 70.70 | 78.61 | 72.61 |
| ZS + definition (E) | 74.15 | 66.24 | 78.00 | 71.97 |
| FS8 GV-domain (E) | 73.83 | 75.16 | 77.84 | 73.89 |
| FS8 cross-domain (E) | 75.68 | 71.34 | 82.49 | 80.25 |
Fine-tuned FLAN-T5
Using the GVFC we have trained FLAN-T5 for framing detection888To fine-tune FLAN-T5, the GVFC was split in train (70%) - dev (10%) - test (20%). Hyperparameters are described in Appendix B. It is important to note that in the experiments of Table 1 we use the entire GVFC dataset as the test set, whereas in this experiment the GVFC test set is considerably smaller to enable fine-tuning.: after task-specific fine-tuning, FLAN-T5 Large has reached an accuracy of 78.83% on the GVFC test set, performing better than the GPT models, as shown in table 6. The table also shows the results of FLAN-T5 Large’s performance on the ITW dataset, compared to the out-of-the-box GPT models: while fine-tuned FLAN-T5 outperforms both GPT models on the GVFC test set, it falls behind on a more varied dataset that does not contain headlines related to the issue of Gun Violence, reaching an accuracy of 68.78% compared to GPT-4’s accuracy of 80.25%.
| GPT-3.5-Turbo | GPT-4 | FLAN-T5 Large | |
| fine-tuned | |||
| Acc. | Acc. | Acc. | |
| GVFC | 65.06 | 67.18 | 78.73 |
| ITW | 71.34 | 80.25 | 68.78 |
6 Conclusions
This work advances our understanding of the performance of pre-trained models, including GPT-4, GPT-3.5 Turbo, and Flan-T5 (small, base, and large), in detecting framing within news headlines without the need for additional fine-tuning, given the scarcity of annotated data for this task. Through a comprehensive experimental approach we found that GPT-4 excels in few-shot scenarios and when asked to explain its results. However, it also showed a tendency to classify headlines as framed when they contain emotional language, while GPT-3.5 showed more resilience against such biases.
Our results indicate that the models are sensitive to the wording of the prompt and the selected examples. However, including a linguistic definition of framing and prompting the model to explain its results leads to more reliable performances. We verify this finding by collecting a new dataset that better resembles real-world use cases for framing detection, where the test data encompasses various topics and the topics are not known in advance.
Furthermore, our study also identifies a performance shortfall in Flan-T5, underscoring the challenges smaller models face in complex detection tasks without specific fine-tuning. As we have shown, after fine-tuning, FLAN-T5’s performance increases surpassing GPT models on in-domain data. However, the situation is reversed again when facing more diverse data (ITW), suggesting that, when the domain is not known in advance, fine-tuning might not be the best choice.
Finally, through our analysis of the GVFC annotations and clear/contested subsets, we noticed the potential of these models in flagging annotation inaccuracies in new or existing datasets, when the predictions are consistent across multiple models. This would be particularly effective when checking annotations, if there is strong agreement between models on a label that contrasts the gold label.
7 Limitations
The findings of this study have to be seen in light of some limitations. For example, our evaluation focuses solely on English-language content, leaving space for further investigation on other languages to explore our findings’ applicability to non-English contexts. This limitation suggests a need for further investigation into the performance of LLMs across different languages and cultural contexts to fully assess the potential use of these models in social science research for detecting framing and analysing media narratives.
Appendix
Appendix A FLAN-T5 Performance Lag
Table 7 shows the results of FLAN-T5 without any specific fine-tuning.
@@ Appendix A FLAN-T5 Performance Lag
| F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | F1 | Acc. | ||
| ZS | No Definition | 0.46 | 49.79 | 17.96 | 49.21 | 0 | 50.10 | 25.86 | 52.74 | 34.33 | 46.19 | 51.01 | 49.06 |
| +Definition | 0.15 | 50.13 | 1.21 | 50.10 | 0 | 50.10 | 0.61 | 50.06 | 42.27 | 48.06 | 37.63 | 43.96 | |
| FS | 2 examples | 0.31 | 50.06 | 0.31 | 50.06 | 0 | 50.10 | 3.12 | 49.98 | 5.74 | 49.64 | 8.32 | 49.64 |
| 8 (focused in-domain) | 0.15 | 50.17 | 3.84 | 50.10 | 0 | 50.13 | 4.83 | 50.17 | 16.78 | 47.53 | 36.85 | 44.69 | |
| 8 (varied in-domain) | 0 | 50.13 | 0.15 | 49.94 | 0 | 50.13 | 1.66 | 49.94 | 22.78 | 45.96 | 35.03 | 45.27 | |
| 8 (cross domain) | 5.30 | 49.37 | 2.95 | 49.60 | 0 | 50.10 | 4.99 | 50.33 | 40.23 | 45.11 | 51.64 | 46.88 | |
| 8 (mixed domain) | 0.31 | 50.21 | 1.51 | 50.02 | 0 | 50.10 | 3.84 | 50.13 | 34.39 | 44.58 | 44.05 | 43.43 | |
Appendix B Detailed Evaluation Fine-tuned FLAN-T5 and Hyperparameters
Table 8 presents the detailed performance comparison on the GVFC Test split and on the ITW dataset; the models are out-of-the-box GPT-3.5-Turbo and GPT-4, and fine-tuned FLAN-T5 Large.
Table 9 presents the detailed performance comparison of fine-tuned FLAN-T5 small, base and large on the GVFC test set.
The FLAN-T5 models (small, base, and large) were fine-tuned using the following hyperparameters: All models employ a learning rate of 5e-05, with a maximum input length of 70 tokens, a maximum label length of 4 tokens, and a batch size of 64. The number of epochs varies with 10 for the small, 40 for the base, and 17 for the large model, each within a 50-epoch limit and with early stopping set at 5 epochs to prevent overfitting.
We used one A100 80G GPU, requiring less than 15 minutes of GPU time per model.
| GPT-3.5-Turbo | GPT-4 | FLAN-T5 Large | ||||||||||
| fine-tuned | ||||||||||||
| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | |
| GVFC | 67.39 | 65.06 | 60.52 | 76.02 | 69.08 | 67.18 | 62.42 | 77.33 | 76.33 | 78.83 | 80.63 | 72.46 |
| ITW | 75.68 | 71.34 | 68.63 | 84.34 | 82.49 | 80.25 | 77.66 | 87.95 | 69.18 | 68.78 | 72.36 | 66.26 |
| FLAN-T5 Small | FLAN-T5 Base | FLAN-T5 Large | ||||||||||
| fine-tuned | fine-tuned | fine-tuned | ||||||||||
| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | |
| GVFC | 0 | 52.29 | 0 | 0 | 79.77 | 79.50 | 74.82 | 85.42 | 76.33 | 78.83 | 80.63 | 72.46 |
| ITW | 2.38 | 47.70 | 100 | 1.20 | 63.87 | 56.05 | 56.48 | 73.49 | 69.18 | 68.78 | 72.36 | 66.26 |
Appendix C Impact of Emotional Language on Framing Detection
Table 10 shows some examples of the impact of emotional language on framing detection.
@@ Appendix C Impact of Emotional Language on Framing Detection
| Headline: Devastated’ Tamera Mowry Confirms Niece Alaina Housley Died In Thousand Oaks Shooting |
@@
| GPT-4 Explanation: The claim is framed as it emphasizes the emotional state of Tamera Mowry (being ’devastated’) while confirming the death of her niece |
| in the shooting. This can elicit an emotional response from the audience and influence their perception of the issue. |
| GPT-3.5 Explanation: The claim is not framed because it is simply reporting the confirmation of Tamera Mowry about the death of her niece in the Thousand |
| Oaks shooting.It does not select any specific aspects or favor a certain interpretation of the issue. |
@@
Appendix D Detailed Results GPT Evaluation
Tables 11 and 12 provide the detailed version of the results of the GPT Evaluation, including Precision and Recall.
| GPT-3.5 Turbo | |||||||||
| Explainable | |||||||||
| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | ||
| Zero-Shot | No Definition | 20.48 | 52.70 | 63.71 | 12.20 | 64.96 | 58.40 | 52.25 | 74.52 |
| +Definition | 51.55 | 59.14 | 63.16 | 43.55 | 65.36 | 63.84 | 58.60 | 59.46 | |
| Few-Shot | 2 examples | 26.83 | 54.16 | 66.06 | 16.83 | 58.25 | 65.84 | 59.32 | 73.98 |
| 8 (focused in-domain) | 40.72 | 58.25 | 69.62 | 28.73 | 64.50 | 63.30 | 59.43 | 72.51 | |
| 8 (varied in-domain) | 46.36 | 60.22 | 70.91 | 34.44 | 68.51 | 65.38 | 58.90 | 71.04 | |
| 8 (cross domain) | 41.22 | 57.67 | 67.19 | 29.73 | 59.42 | 66.04 | 62.79 | 58.86 | |
| 8 (mixed domain) | 56.93 | 63.49 | 69.25 | 48.34 | 63.33 | 64.37 | 61.82 | 67.64 | |
| GPT-4 | |||||||||
| Explainable | |||||||||
| F1 | Acc. | Prec. | Rec. | F1 | Acc. | Prec. | Rec. | ||
| Zero-Shot | No Definition | 64.96 | 58.40 | 56.05 | 77.22 | 63.75 | 60.41 | 58.71 | 69.73 |
| +Definition | 65.36 | 63.84 | 62.63 | 68.34 | 64.64 | 60.99 | 59.03 | 71.43 | |
| Few-Shot | 2 examples | 58.25 | 65.84 | 74.73 | 47.72 | 64.23 | 64.92 | 65.41 | 63.09 |
| 8 (focused in-domain) | 64.50 | 63.30 | 62.36 | 66.80 | 66.44 | 63.96 | 60.06 | 71.51 | |
| 8 (varied in-domain) | 68.51 | 65.38 | 62.75 | 75.44 | 70.41 | 66.92 | 63.61 | 78.84 | |
| 8 (cross domain) | 59.42 | 66.04 | 73.63 | 49.81 | 59.09 | 64.61 | 69.86 | 51.20 | |
| 8 (mixed domain) | 63.33 | 64.37 | 65.14 | 61.62 | 63.87 | 63.76 | 63.58 | 64.17 | |
Appendix E Detailed Results In The Wild Evaluation
Table 13 shows the detailed results of the evaluation In The Wild, including Precision and Recall.
| GPT-3.5 Turbo | GPT-4 | |||||||
| F1 | Acc. | Prec | Rec. | F1 | Acc. | Prec | Rec. | |
| ZS + definition (S) | 74.16 | 70.70 | 69.47 | 79.52 | 78.61 | 72.61 | 66.95 | 95.18 |
| ZS + definition (E) | 74.15 | 66.24 | 62.30 | 91.57 | 78.00 | 71.97 | 66.67 | 93.98 |
| FS8 GV-domain | 73.83 | 75.16 | 83.33 | 66.27 | 77.84 | 73.84 | 70.59 | 86.75 |
| FS8 cross-domain | 75.68 | 71.34 | 68.63 | 84.34 | 82.49 | 80.25 | 77.66 | 87.95 |